OpenStudy (samara8954):

A recipe calls for of a cup of milk for 12 cookies. How many cups of milk are needed to make 180 cookies? 6 12 5 60

OpenStudy (samara8954):

1/3

OpenStudy (samara8954):

@calculusxy

OpenStudy (aripotta):

make a proportion\[\frac{ \frac{ 1 }{ 3 } }{ 12 }=~\frac{ x }{ 180 }\] cross-ultiply and solve for x

OpenStudy (aripotta):

cross-multiply*

OpenStudy (samara8954):

so the answer is

OpenStudy (aripotta):

to cross-multiply, you multiply the first fraction's numerator with the second fraction's denominator. and multiply the first fraction's denominator with the second fraction's numerator

OpenStudy (samara8954):

what is the answer

OpenStudy (aripotta):

 so 1/3(180) = 12x